# Minkowski tensors for voxelized data: robust, asymptotically unbiased estimators

### v1

@@ 4 Algorithms to estimate Minkowski tensors
In this section, we define explicit algorithms that are used to estimate the Minkowski tensors, given a discrete (pixelated) data set. We focus on the current improvements in comparison to the original algorithm suggested in [21]. The digitized estimator of the volume tensors and especially the least squares fit with respect to the radial dependence strongly improve the robustness of the algorithm. Then we discuss details of our open-source implementation [47] and provide some advice on the choice of parameters. Finally, we demonstrate the performance for some test cases; see Sections 4.4 and 4.5.
4.1 Description and analysis
In the following, we describe two methods for estimating Minkowski tensors. The first is based on Corollary 4.1 and works for arbitrary finite unions of compact sets with positive reach and (surface) tensor functionals $\Phi^{r,s}_{d-1}$ with $s\geq 1$. The second approach, which is applicable for arbitrary tensor functionals $\Phi^{r,s}_{j}$, is based on solving a least squares problem and offers practical advantages. However, its theoretical foundation is only available for sets with positive reach. Nevertheless, in Section 4.5, we will see that it provides surprisingly accurate results for a set without positive reach.
To explain the two approaches, we fix integers $r,s\in{\mathbb{N}}_{0}$. The Voronoi tensor $\mathcal{V}_{R}^{r,s}(K)$ with distance parameter $R\geq 0$ and rank parameter tuple $(r,s)\in{\mathbb{N}}_{0}^{2}$ of a nonempty compact set $K\subset{\mathbb{R}}^{d}$ is defined by
@@ 4.1 Description and analysis
where $K^{R}$ is the set of points with distance at most $R>0$ from $K$ and $p_{K}$ is the (almost everywhere uniquely defined) metric projection on $K$. For $r=s=0$ the Voronoi tensor with distance parameter $R$ is the volume of $K^{R}$, i.e., the volume of the set of points with distance at most $R$ from $K$.
The following corollary of Theorem 3.4 shows how surface tensors can be approximated by Voronoi tensors. Recall the definition of the reach measure $\mu_{d-1}(A;\cdot)$ of a nonempty compact subset of ${\mathbb{R}}^{d}$ from Section 3, which equals the support measure $\Lambda_{d-1}(A,\cdot)$ if $A\in\mathcal{U}^{d}$ (see Remark 3.1 which ensures that the Minkowski tensor $\Phi^{r,s}_{d-1}$ in (4.2) is consistently defined for all $r,s\in{\mathbb{N}}_{0}$). An explicit description of the reach measure $\mu_{d-1}(A;\cdot)$ is provided in (3.4) (see [23, Proposition 4.1]).
Corollary 4.1.
Let $K\subset{\mathbb{R}}^{d}$ be a finite union of compact sets of positive reach. If $r\in{\mathbb{N}}_{0}$ and $s\in{\mathbb{N}}$, then
$\lim_{\varepsilon\to 0_{+}}\frac{1}{\varepsilon^{1+s}}\mathcal{V}^{r,s}_{\varepsilon}(K)=r!s!\kappa_{s+1}\Phi^{r,s}_{d-1}(K)$
and
$\lim_{\varepsilon\to 0_{+}}\frac{1}{\varepsilon}\left(\mathcal{V}^{r,0}_{\varepsilon}(K)-\mathcal{V}^{r,0}_{\varepsilon^{2}}(K)\right)=r!2\,\Phi^{r,0}_{d-1}(K),$
where, for $r,s\in{\mathbb{N}}_{0}$,
$\Phi^{r,s}_{d-1}(K):=\frac{1}{r!s!}\frac{2}{\omega_{1+s}}\int_{{\mathbb{R}}^{d}\times\mathbb{S}^{d-1}}x^{r}u^{s}\,\mu_{d-1}(K;\mathrm{d}(x,u)).$ (4.2)
Proof.
$\frac{2}{s+1}\int x^{r}u^{s}\,\mu_{d-1}(K;\mathrm{d}{(}x,u)),$
If $s=0$, then an application of Theorem 3.3 and relation (3.4) yield
$=2r!\Phi^{r,0}_{d-1}(K)-0,$
which proves the assertion. ∎
For a finite set $K_{0}\subset{\mathbb{R}}^{d}$ and $r,s\in{\mathbb{N}}_{0}$, the Voronoi tensors simplify to
@@
where $B(x,R)$ is the closed ball with center $x$ and radius $R$ and
@@
is the Voronoi cell of $x\in K_{0}$ with respect to the set $K_{0}$. It is clear that each compact set $K\subset{\mathbb{R}}^{d}$ can be approximated with any specified precision by finite sets $K_{0}\subset{\mathbb{R}}^{d}$ with respect to the Hausdorff metric $d_{H}$ on compact subsets of ${\mathbb{R}}^{d}$ (see, e.g., (4.10)). If $K,K_{0}\subseteq B(o,\rho)$ for some $\rho>0$, then
@@
which follows from [21, Thm. 4.1], where $C(d,R,\rho,r,s)$ is a constant depending only on the arguments in brackets. An inspection of the proof of Theorem 4.1 in [21] shows that the constant $C(d,R,\rho,r,s)$ can be chosen such that $C(d,R,\rho,r,s)\leq C(d,R_{0},\rho,r,s)$ if $0<R\leq R_{0}$. Note that (4.4) does not require that $K_{0}$ is a subset of $K$ (although an approximation of $K$ from inside can be guaranteed for a given $K$) and holds whenever $K,K_{0}$ are compact subsets of $B(o,\rho)$. Recall from [21] that $d_{H}(K,K_{0})^{\frac{1}{2}}$ can be replaced by $d_{H}(K,K_{0})$ if $r=s=0$.
@@
If the indicator functions are nonzero, then for every $i\in{\mathbb{N}}$ there is some $w\in K_{0}$ such that $\|z_{i}+U_{a}-w\|\leq R$, and hence $z_{i}\in K_{0}+B(\sqrt{d}a+R)$. In particular, this shows that the effective ranges of the summations over $i\geq 1$ and $w\in K_{0}$ are finite and can be chosen independently of $U_{a}$. Moreover, (2.2) yields
@@
For this, observe that almost surely $p_{K_{0}}(z+U_{a})$ is uniquely defined for each $z\in a\cdot{\mathbb{Z}}^{d}$. Hence, almost surely
@@
Note that in the last step integration and summation can be interchanged, since for each $x\in K_{0}$ there are only finitely many summands $z\in a\cdot{\mathbb{Z}}^{d}$ such that the first indicator is nonzero. Moreover, we have
@@
×$\qquad\qquad\qquad\qquad\times\mathds{1}_{z+[0,a)^{d}}(y)\,\mathcal{H}^{d}(\mathrm{d}y).$
If $z+[0,a)^{d}\subseteq B(x,R)\cap{V}_{x}(K_{0})$, then (2.2) and (2.3) shows that the expression ||$\cdot$|| ⋅ | under the integral can be bounded from above by
@@
If $(z+[0,a)^{d})\cap B(x,R)\cap{V}_{x}(K_{0})=\emptyset$, then the expression is zero. In the remaining case, we have $(z+[0,a)^{d})\cap\partial(B(x,R)\cap{V}_{x}(K_{0}))\neq\emptyset$ and the expression ||$\cdot$|| ⋅ | is bounded from above by $2R^{s}$ (see (2.2)). Thus we obtain
@@
$\leq s\cdot d\rho^{r}R^{s-1}\cdot a\cdot\mathcal{H}^{d}\left(K+B(o,d_{H}(K,K_{0})+R)\right)$
@@
where we used that the Voronoi cells have pairwise disjoint interiors. Since $K,K_{0}\subset B(o,\rho)$, we get
$\mathcal{H}^{d}(K+B(o,d_{H}(K,K_{0})+R))\leq\mathcal{H}^{d}\left(B(o,2\rho+R)\right)\leq\kappa_{d}(2\rho+R)^{d}.$ (4.8)
@@
$=\left[\left(1+\frac{\sqrt{d}\cdot a}{r_{x}\wedge R}\right)^{d}-1\right]\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right)$
@@
where we use that $\operatorname{{\sf N}}(K_{0})\leq 2\rho$ if $R<\operatorname{{\sf N}}(K_{0})$. If $R\geq\operatorname{{\sf N}}(K_{0})$, then we can choose $c_{2}(d,\rho,R)=c_{1}(d)$. The sum of the volumes $\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right)$ over $x\in K_{0}$ is bounded from above by $\kappa_{d}(\rho+R)^{d}$. Combination of (4.1), (4.8), and (4.1) finally yields
@@
$\leq d\rho^{r}R^{s-1}(2\rho+R)^{d}\cdot\left(s\cdot a+R\cdot c_{2}(d,\rho,R)\cdot\max\left\{\frac{a}{\operatorname{{\sf N}}(K_{0})},\left(\frac{a}{\operatorname{{\sf N}}(K_{0})}\right)^{d}\right\}\right),$
@@
where $C^{\prime\prime}(d,R_{0},\rho,r,s)>0$ is a constant depending only on $d,R_{0},\rho,r,s$, and $d_{H}(K,K_{0})^{\frac{1}{2}}$ can be replaced by $d_{H}(K,K_{0})$ if $r=s=0$.
Combination of Corollary 4.1 and Corollary 4.4 finally yields the next result. It serves as a theoretical foundation for the approximation of surface tensors of finite unions of sets with positive reach via Corollary 4.1. We remark that for any nonempty compact set $K\subset{\mathbb{R}}^{d}$, the finite set
@@
Let $K\subset{\mathbb{R}}^{d}$ be a finite union of compact sets with positive reach, and let $r\in{\mathbb{N}}_{0}$, $s\in{\mathbb{N}}$. Let $\rho>0$ be such that $K\subseteq B(o,\rho)$. If $(\varepsilon_{n})_{n\in{\mathbb{N}}}$ satisfies $\varepsilon_{n}\to 0_{+}$ as $n\to\infty$, then there exist sequences $(K_{0}(n))_{n\in{\mathbb{N}}}$ of finite subsets of ${\mathbb{R}}^{d}$ with $K_{0}(n)\to K$ as $n\to\infty$ and $(a_{n})_{n\in{\mathbb{N}}}$ of real numbers such that, almost surely,
@@
We can assume that $\varepsilon_{m}\leq\varepsilon(1)$ for $m\in{\mathbb{N}}$. For $m\in{\mathbb{N}}$ there is a unique $n\in{\mathbb{N}}$ such that $\varepsilon_{m}\in(\varepsilon(n+1),\varepsilon(n)]$. By the remarks preceding Theorem 4.5 , there is a finite set $K_{0}(m)\subset{\mathbb{R}}^{d}$ such that
@@
$0<a_{m}\leq\operatorname{{\sf N}}(K_{0}(m))\wedge\varepsilon_{m}^{2}\leq\operatorname{{\sf N}}(K_{0}(m))\wedge\varepsilon_{m}\leq 1$
@@
$\left|\frac{1}{\varepsilon_{m}}\left(\widehat{\mathcal{V}}_{\varepsilon_{m}}^{r,0}(K_{0}(m),a_{m})-\widehat{\mathcal{V}}_{\varepsilon_{m}^{2}}^{r,0}(K_{0}(m),a_{m})\right)-r!2\,\Phi^{r,0}_{d-1}(K)\right|\leq\frac{1}{n},$
@@
Using the estimator from (4.6), Corollary 4.1 can in principle be applied directly to estimate the tensors $\Phi^{r,s}_{d-1}(K)$ from Corollary 4.1 by choosing a very small value for $R$, whenever $K$ is a finite union of compact sets with positive reach and $s\geq 1$. We tested the performance for a rectangle $[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$ (see Example 4.9), a spherical shell with inner radius 1 and outer radius 2 (see Example 4.10), and for a rectangle from which a smaller (open) rectangle has been removed (see Example 4.11). The generated finite test data are obtained by intersecting these objects with a grid. The simulated estimates can be found in Table 2. To achieve good results, however, we needed an $a$ that is significantly smaller than $R$, which itself must already be small. Since a small value of $a$ considerably slows down the computation of the estimator from (4.6), this is a substantial disadvantage in practice. Therefore, for estimating the tensors from Corollary 4.1 we propose to use the alternative approach via a least squares problem, whenever it is applicable. For example, a single rendition (i.e., a run of the algorithm that provides an estimate) for the spherical shell $S$ in the simulations for Table 2 took about 21 minutes on a standard personal computer, whereas in the simulations from Table 5 with the approach via a least squares problem, all 25 renditions for the same spherical shell took only about 3 minutes on a standard personal computer. Note that two renditions differ in the grid (4.5) used, as randomness only comes into play here. Surprisingly, the alternative approach led to excellent results even in cases where we only have a heuristic justification (see Table 6 for Example 4.11). Finally, we point out that we observed that the second part of Theorem 4.5, namely the approach for $s=0$, requires far too much runtime and memory capacity for practical applications. Since $a$ must be significantly smaller than $\varepsilon$ and $\varepsilon^{2}$ for the application of the Theorem, and $\varepsilon$ itself should already be a very small value, $a$ becomes too small (and thus the number of points in the lattice process too large) for practical use. However, for $r=s=0$ this problem can be overcome, as the following remark demonstrates.
@@
Let $K$ be a finite union of compact sets with positive reach and $\operatorname{tr}(T)$ the trace of a tensor $T\in\mathbb{T}^{2}({\mathbb{R}}^{d})$, that is, $\operatorname{tr}(T):=\sum_{j=1}^{d}T(e_{j},e_{j})$. Then it is true that

### v2

@@ 4 Algorithms to estimate Minkowski tensors
Based on our theoretical results from the previous section, we now define two explicit algorithms to estimate the Minkowski tensors of a set $K$ that is represented by a point cloud $K_{0}$, e.g., by a pixelated image. For both algorithms, it is necessary to estimate the so-called Voronoi tensors, which is the subject of Subsection 4.1. In the next two Subsections 4.2 and 4.3, we introduce and describe the two algorithms in detail. Finally, in Subsections 4.4 and 4.5, we discuss important aspects of our open-source implementation [60] and provide some advice on the choice of parameters.
4.1 Voronoi tensor estimation
The Voronoi tensor $\mathcal{V}_{R}^{r,s}(K)$ with distance parameter $R\geq 0$ and rank parameter tuple $(r,s)\in{\mathbb{N}}_{0}^{2}$ of a nonempty compact set $K\subset{\mathbb{R}}^{d}$ is defined by
@@ 4.1 Voronoi tensor estimation
where $K^{R}$ is the set of points with distance at most $R>0$ from $K$ and $p_{K}$ is the (almost everywhere uniquely defined) metric projection on $K$. The Voronoi tensor $\mathcal{V}_{R}^{r,s}(K)$ is the total Voronoi tensor measure $\mathcal{V}_{R}^{r,s}(K;{\mathbb{R}}^{d})$ introduced in [32]. The present work is inspired by the results in [32], which in turn were motivated and based on the contributions [16, 56]. Extensions to $\delta$-Voronoi tensors, where $\delta$ is a distance-like function as considered in [22], which have been introduced to improve the robustness of estimation methods against outliers for the Voronoi covariance measure, will not be pursued here. If $s>0$, then the integral in (4.1) of the symmetric tensor product of the symmetric tensors $p_{K}(x)^{r}$ and $(x-p_{K}(x))^{s}$ can be restricted to $K^{R}\setminus K$ without changing the value of the integral. For $r=s=0$ the Voronoi tensor with distance parameter $R$ is the volume of $K^{R}$. For a finite set $K_{0}\subset{\mathbb{R}}^{d}$ and $r,s\in{\mathbb{N}}_{0}$, the Voronoi tensors simplify to
@@
where $B(x,R)$ is the closed ball with center $x$ and radius $R$ (we omit the upper index that indicates the dimension) and
@@
is the Voronoi cell of $x\in K_{0}$ with respect to the set $K_{0}$. It is clear that each compact set $K\subset{\mathbb{R}}^{d}$ can be approximated with any specified precision by finite sets $K_{0}\subset{\mathbb{R}}^{d}$ with respect to the Hausdorff metric $d_{H}$ on compact subsets of ${\mathbb{R}}^{d}$ (see, e.g., (4.14)); we refer to [67, p. 571] for a definition of the Hausdorff metric. If $K,K_{0}\subseteq B(o,\rho)$ for some $\rho>0$, then
@@
which follows from [32, Theorem 4.1], where $C(d,R,\rho,r,s)$ is a constant depending only on the arguments in brackets. An inspection of the proof of Theorem 4.1 in [32] shows that the constant $C(d,R,\rho,r,s)$ can be chosen such that $C(d,R,\rho,r,s)\leq C(d,R_{0},\rho,r,s)$ if $0<R\leq R_{0}$. Note that (4.3) does not require that $K_{0}$ is a subset of $K$ (although an approximation of $K$ from inside can be guaranteed for a given $K$) and holds whenever $K,K_{0}$ are compact subsets of $B(o,\rho)$. Recall from [32] that $d_{H}(K,K_{0})^{\frac{1}{2}}$ can be replaced by $d_{H}(K,K_{0})$ if $r=s=0$.
@@
If the indicator functions are nonzero, then for every $i\in{\mathbb{N}}$ there is some $w\in K_{0}$ such that $\|z_{i}+U_{a}-w\|\leq R$, and hence $z_{i}\in K_{0}+B(o,\sqrt{d}a+R)$. In particular, this shows that the effective ranges of the summations over $i\geq 1$ and $w\in K_{0}$ are finite and can be chosen independently of $U_{a}$. Moreover, (2.2) yields
@@
For this, observe that almost surely $p_{K_{0}}(z+U_{a})$ is uniquely defined for each $z\in a\cdot{\mathbb{Z}}^{d}$. Moreover, since $K_{0}$ is finite and there are only finitely many $z\in a\cdot{\mathbb{Z}}^{d}$ having distance at most $\sqrt{d}aR$ from a point in $K_{0}$, all sums encountered in the following are finite. Hence, almost surely
@@
Note that in the last step integration and summation can be interchanged, since for each $x\in K_{0}$ there are only finitely many summands $z\in a\cdot{\mathbb{Z}}^{d}$ such that the first indicator is nonzero (as pointed out above). Moreover, we have
@@
$\qquad\qquad\qquad\qquad\times\mathds{1}_{z+[0,a)^{d}}(y)\,\mathcal{H}^{d}(\mathrm{d}y).$
Suppose that $y\in z+[0,a)^{d}$. If $z+[0,a)^{d}\subseteq B(x,R)\cap{V}_{x}(K_{0})$, then (2.3) shows that the expression $|\cdot|$ under the integral can be bounded from above by
@@
If $(z+[0,a)^{d})\cap B(x,R)\cap{V}_{x}(K_{0})=\emptyset$, then the expression is zero. In the remaining case, we have $(z+[0,a)^{d})\cap\partial(B(x,R)\cap{V}_{x}(K_{0}))\neq\emptyset$ and the expression $|\cdot|$ is bounded from above by $2R^{s}$ (see (2.2)). Thus we obtain
@@
$\leq s\cdot d\rho^{r}R^{s-1}\cdot a\cdot\mathcal{H}^{d}\left(K_{0}+B(o,R)\right)$
@@
where we used that the Voronoi cells have pairwise disjoint interiors. Since $K_{0}\subset B(o,\rho)$, we get
$\mathcal{H}^{d}(K_{0}+B(o,R))\leq\kappa_{d}(\rho+R)^{d}.$ (4.7)
@@
$=2\left[\left(1+\frac{\sqrt{d}\cdot a}{r_{x}\wedge R}\right)^{d}-1\right]\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right)$
@@
where we can choose $c_{2}(d,\rho,R):=c_{1}(d)$ if $R\geq\operatorname{{\sf N}}(K_{0})$ and
$c_{2}(d,\rho,R):=c_{1}(d)\max\left\{\frac{2\rho}{R},\left(\frac{2\rho}{R}\right)^{d}\right\}$
if $R<\operatorname{{\sf N}}(K_{0})$, since $\operatorname{{\sf N}}(K_{0})\leq 2\rho$. The sum of the volumes $\mathcal{H}^{d}\left(B(x,R)\cap{V}_{x}(K_{0})\right)$ over $x\in K_{0}$ is bounded from above by $\kappa_{d}(\rho+R)^{d}$ (as used before). Combination of (4.1), (4.7), and (4.1) finally yields
@@
$\leq d\rho^{r}R^{s-1}(\rho+R)^{d}\cdot\left(s\cdot a+R\cdot c_{2}(d,\rho,R)\cdot\max\left\{\frac{a}{\operatorname{{\sf N}}(K_{0})},\left(\frac{a}{\operatorname{{\sf N}}(K_{0})}\right)^{d}\right\}\right),$
@@
where $C^{\prime\prime}(d,R_{0},\rho,r,s)>0$ is a constant depending only on $d,R_{0},\rho,r,s$, and $d_{H}(K,K_{0})^{\frac{1}{2}}$ can be replaced by $d_{H}(K,K_{0})$ if $r=s=0$. Moreover, if $0<a\leq\operatorname{{\sf N}}(K_{0})\leq R$, then
$|\widehat{\mathcal{V}}_{R}^{r,s}(K_{0},a)-\mathcal{V}_{R}^{r,s}(K_{0})|\leq d\rho^{r}R^{s}(\rho+R)^{d}(s+c_{1}(d))\,\frac{a}{\operatorname{{\sf N}}(K_{0})}.$ (4.9)
Remark 4.4.
Increasing the number of points in $K_{0}$ admits a better approximation of $K$ by $K_{0}$. This part of the approximation is already covered in previous work, which we summarized in (4.3). The main contribution of Theorem 4.2 is to bound $|\widehat{\mathcal{V}}^{r,s}_{R}(K_{0},a)-{\mathcal{V}}^{r,s}_{R}(K_{0})|$ from above. An explicit upper bound is stated in (4.9). In the derivation of this bound, we encountered two cases. In the first case, we considered $z\in a\cdot{\mathbb{Z}}^{d}$ for which $(z+[0,a)^{d})\subseteq B(x,R)\cap{V}_{x}(K_{0})$. For the corresponding sum, we get the upper bound $d\rho^{r}R^{s-1}(\rho+R)^{d}sa$, which improves as $a$ becomes smaller, independently of $N(K_{0})$. The second main case deals with $z\in a\cdot{\mathbb{Z}}^{d}$ for which $(z+[0,a)^{d})\cap\partial(B(x,R)\cap{V}_{x}(K_{0}))\neq\emptyset$. For the corresponding second sum in (4.1) we obtain the alternative upper bound
$d\rho^{r}R^{s}\sum_{x\in K_{0}}\mathcal{H}^{d}\left(\partial(B(x,R)\cap{V}_{x}(K_{0}))+B(o,\sqrt{d}\cdot a)\right)$
$\leq c_{d}\rho^{r}R^{s}|K_{0}|\cdot a\cdot\sum_{i=1}^{d}R^{d-i}a^{i-1}.$ (4.10)
A simple volume bound shows that $|K_{0}|\leq\left(\frac{2\rho}{N(K_{0})}+1\right)^{d}$, which cannot be improved in general. In this way, we obtain an upper bound that involves ratios $\frac{a}{N(K_{0})^{i}}$ also with $i>1$ instead of just the ratio $\frac{a}{N(K_{0})}$. As a consequence of our approach, increasing $K_{0}$ requires more boundary terms to be controlled. If $K_{0}$ is considered to be fixed, then of course the error term decreases linearly with $a$. In a simulation study, we found that keeping $a$ fixed and increasing $|K_{0}|$ (or decreasing $N(K_{0})$) when approximating a rectangle $K$ by finite point clouds $K_{0}$ does not improve the approximation of ${\mathcal{V}}^{r,s}_{R}(K_{0})$ by $\widehat{\mathcal{V}}^{r,s}_{R}(K_{0},a)$. It seems plausible that increasing $|K_{0}|$ (and decreasing $N(K_{0})$, accordingly) should be matched by a corresponding decrease of the grid size $a$.
Remark 4.5.
The task of finding a useful upper bound for the volume of all points having distance at most $\sqrt{d}\cdot a$ from $\partial(B(x,R)\cap{V}_{x}(K_{0}))$ in (4.1) is reminiscent of the situation treated in [49, Theorem 3]. However, the assumption that the boundary of the set $X$ there should have positive reach is clearly not satisfied by the boundary of the Voronoi cell $V_{x}(K_{0})$.
4.2 Voronoi-FD algorithm
In the following, we describe a first method for estimating Minkowski tensors, which we call the Voronoi-FD algorithm, where “FD” stands for finite difference. The method is based on Corollary 4.6 and works for arbitrary finite unions of compact sets with positive reach and (surface) tensor functionals $\Phi^{r,s}_{d-1}$ with $s\geq 1$.
The following corollary of Theorem 3.9 shows how surface tensors can be approximated by Voronoi tensors. Recall the definition of the reach measure $\mu_{d-1}(A;\cdot)$ of a nonempty compact subset of ${\mathbb{R}}^{d}$ from Section 3, which equals the support measure $\Lambda_{d-1}(A,\cdot)$ if $A\in\mathcal{U}^{d}$ (see Remark 3.2 which ensures that the Minkowski tensor $\Phi^{r,s}_{d-1}$ in (4.13) is consistently defined for all $r,s\in{\mathbb{N}}_{0}$). An explicit description of the reach measure $\mu_{d-1}(A;\cdot)$ is provided in (3.5) (see [28, Proposition 4.1]).
Corollary 4.6.
Let $K\subset{\mathbb{R}}^{d}$ be a nonempty compact set such that (3.3) holds. If $r\in{\mathbb{N}}_{0}$ and $s\in{\mathbb{N}}$, then
$\lim_{\varepsilon\to 0_{+}}\frac{1}{\varepsilon^{1+s}}\mathcal{V}^{r,s}_{\varepsilon}(K)=r!s!\kappa_{s+1}\Phi^{r,s}_{d-1}(K)$ (4.11)
and
$\lim_{\varepsilon\to 0_{+}}\frac{1}{\varepsilon}\left(\mathcal{V}^{r,0}_{\varepsilon}(K)-\mathcal{V}^{r,0}_{\varepsilon^{2}}(K)\right)=r!2\,\Phi^{r,0}_{d-1}(K),$ (4.12)
where, for $r,s\in{\mathbb{N}}_{0}$,
$\Phi^{r,s}_{d-1}(K):=\frac{1}{r!s!}\frac{2}{\omega_{1+s}}\int_{{\mathbb{R}}^{d}\times\mathbb{S}^{d-1}}x^{r}u^{s}\,\mu_{d-1}(K;\mathrm{d}(x,u))$ (4.13)
and $\mu_{d-1}(K;\cdot)$ is a nonnegative Borel measure that satisfies (3.5).
Proof.
$\frac{2}{s+1}\int x^{r}u^{s}\,\mu_{d-1}(K;\mathrm{d}{(}x,u)),$
If $s=0$, then an application of Theorem 3.7 and relation (3.5) yield
$=2r!\Phi^{r,0}_{d-1}(K),$
which proves the assertion. ∎
Remark 4.7.
Corollary 4.6 applies, for instance, when $A$ is a finite union of compact sets of positive reach (by Lemma 3.4) or $A$ is a parallel set of an arbitrary compact set (by Remark 3.5). The discussion in Remark 3.8 shows that no general statement is possible concerning the speed of convergence in (4.11) or (4.12).
Combination of Corollary 4.3 and Corollary 4.6 finally yields the next result. It serves as a theoretical foundation for the approximation of surface tensors of finite unions of sets with positive reach via Corollary 4.6. We remark that for any nonempty compact set $K\subset{\mathbb{R}}^{d}$, we can approximate it to arbitrary precision with a discrete set, for example, the finite set
@@ 4.2 Voronoi-FD algorithm
Let $K\subset{\mathbb{R}}^{d}$ be a nonempty compact set such that (3.3) holds. Let $r\in{\mathbb{N}}_{0}$, $s\in{\mathbb{N}}$, and let $\rho>0$ be such that $K\subseteq B(o,\rho)$. If $(\varepsilon_{n})_{n\in{\mathbb{N}}}$ satisfies $\varepsilon_{n}\to 0_{+}$ as $n\to\infty$, then there exist sequences $(K_{0}(n))_{n\in{\mathbb{N}}}$ of finite subsets of ${\mathbb{R}}^{d}$ with $K_{0}(n)\to K$ in the Hausdorff distance as $n\to\infty$ and $(a_{n})_{n\in{\mathbb{N}}}$ of real numbers such that, almost surely,
@@
We can assume that $\varepsilon_{m}\leq\varepsilon(1)$ for $m\in{\mathbb{N}}$. For $m\in{\mathbb{N}}$ there is a unique $n\in{\mathbb{N}}$ such that $\varepsilon_{m}\in(\varepsilon(n+1),\varepsilon(n)]$. By the remarks preceding Theorem 4.8, there is a finite set $K_{0}(m)\subset B(o,\rho)$ (if $m$ and therefore $n$ is sufficiently large) such that
@@
$0<a_{m}\leq\operatorname{{\sf N}}(K_{0}(m))\wedge\varepsilon_{m}^{2}\leq\operatorname{{\sf N}}(K_{0}(m))\wedge\varepsilon_{m}\leq 1.$
@@
$\left|\frac{1}{\varepsilon_{m}}\left(\widehat{\mathcal{V}}_{\varepsilon_{m}}^{r,0}(K_{0}(m),a_{m})-\widehat{\mathcal{V}}_{\varepsilon_{m}^{2}}^{r,0}(K_{0}(m),a_{m})\right)-r!2\,\Phi^{r,0}_{d-1}(K)\right|\leq\frac{1}{n},$ (4.19)
@@
Suppose that $\varepsilon_{n}=\varepsilon(n)$ for $n\in{\mathbb{N}}$, where $\varepsilon(n)$ is chosen so that (4.15) is valid if $s=0$ and
$\left|{\varepsilon^{-(1+s)}}\mathcal{V}^{r,s}_{\varepsilon}(K)-r!s!\kappa_{s+1}\,\Phi^{r,s}_{d-1}(K)\right|\leq\frac{1}{2n}\quad\text{if }0<\varepsilon\leq\varepsilon(n)$ (4.20)
if $s\in{\mathbb{N}}$. An inspection of the proof of Theorem 4.8 shows that, for fixed $d,\rho,r,s,R$, we can choose $a_{n}$ of the order $n^{-2}\varepsilon_{n}^{3(1+s)}$ so that (4.19) holds (with $m=n$) for $s=0$ and, similarly,
$\left|\varepsilon_{n}^{-(1+s)}\,\widehat{\mathcal{V}}_{\varepsilon_{n}}^{r,s}(K_{0}(n),a_{n})-r!s!\kappa_{s+1}\,\Phi^{r,s}_{d-1}(K)\right|\leq\frac{1}{n}$
if $s\in{\mathbb{N}}$.
An explicit choice of the sequence $\varepsilon(n)$, $n\in{\mathbb{N}}$, in (4.15) and (4.20) will have to depend on the underlying set $K$, as the examples in Remark 3.8 demonstrate (see also Remark 4.7). In the special case where $K$ has positive reach, we can choose $\varepsilon_{n}$ of the order $n^{-1}$ and $a_{n}$ of the order $n^{-5-3s}$.
The Voronoi-FD algorithm estimates the tensors $\Phi^{r,s}_{d-1}(K)$ from Corollary 4.6, where $K$ is a nonempty compact set such that (3.3) holds, by using (4.11) if $s\geq 1$ and (4.12) if $s=0$. In particular, the algorithm can be applied for a set $K$ that is a finite union of compact sets with positive reach. A very small value for $\varepsilon$ is chosen for this purpose, and the estimator from (4.5) is used. We tested the performance for a rectangle $\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$ (see Example 5.1), a spherical shell with inner radius 1 and outer radius 2 (see Example 5.2), and for a rectangle from which a smaller (open) rectangle has been removed (see Example 5.3). The generated finite test data are obtained by intersecting these objects with a grid. The simulated estimates can be found in Table 2. To achieve good results, however, we needed an $a$ that is significantly smaller than $\varepsilon$, which itself must already be small (compare with (4.9), which guarantees a good approximation if $a\ll\operatorname{{\sf N}}(K_{0})\leq\varepsilon=R$). In practice, the choice of $\varepsilon$ can be limited by the available $av(K_{0})$, defined in (4.23), and the typical length scale of the spatial structure. Since a small value of $a$ considerably slows down the computation of the estimator from (4.5), this is a substantial disadvantage in practice. Therefore, for estimating the tensors from Corollary 4.6 we propose to use the alternative approach via a least squares problem (Voronoi-LSQ algorithm) described in Subsection 4.3, whenever it is applicable. For example, a single rendition (i.e., a run of the algorithm that provides an estimate) for the spherical shell $S$ in the simulations for Table 2 using the Voronoi-FD algorithm took about 21 minutes on a standard personal computer, whereas in the simulations from Table 5 using the Voronoi-LSQ algorithm, all 25 renditions for the same spherical shell took only about 3 minutes on a standard personal computer. Note that two renditions differ in the stationarized grid (4.4) since a different random offset of the grid is chosen for each rendition. Surprisingly, the Voronoi-LSQ algorithm led to excellent results even in cases where we only have a heuristic justification (see Table 6 for Example 5.3). Finally, we point out that we observed that the second part of Theorem 4.8, namely the approach for $s=0$, requires far too much runtime and memory capacity for practical applications. Since $a$ must be significantly smaller than $\varepsilon$ and $\varepsilon^{2}$ for the application of the Theorem, and $\varepsilon$ itself should already be a very small value, $a$ becomes too small (and thus the number of points in the lattice process too large) for practical use. However, for $r=s=0$ this problem can be overcome, as the following remark demonstrates.
Remark 4.10.
Let $K$ be a nonempty compact set such that (3.3) holds (which is satisfied if $K$ is a finite union of compact sets with positive reach). If $\operatorname{tr}(T)$ denotes the trace of a tensor $T\in\mathbb{T}^{2}({\mathbb{R}}^{d})$, that is, $\operatorname{tr}(T):=\sum_{j=1}^{d}T(e_{j},e_{j})$, then